# The inverse initial data problem for anisotropic Navier–Stokes equations via Legendre time reduction method

### v2

@@ 5 Numerical study
In this section, we describe a semi-implicit finite difference scheme to generate simulated data, and the Picard quasi-reversibility method employed to solve the time-reduced model (3.17), thereby addressing Problem 1. We also present several numerical examples to illustrate the performance of the proposed approach.
5.1 Data generation
We present numerical simulations in the two-dimensional setting ($d=2$) for simplicity. The fourth-order viscosity tensor $\bm{\mu}$ used in our simulations is specified via its flattened representation:
$\bm{\mu}_{\rm flat}=\begin{bmatrix}\mu_{1111}&\mu_{1112}&\mu_{1121}&\mu_{1122}\\ \mu_{1211}&\mu_{1212}&\mu_{1221}&\mu_{1222}\\ \mu_{2111}&\mu_{2112}&\mu_{2121}&\mu_{2122}\\ \mu_{2211}&\mu_{2212}&\mu_{2221}&\mu_{2222}\end{bmatrix}=\begin{bmatrix}1&\tfrac{1}{16}&\tfrac{1}{16}&\tfrac{3}{8}\\ \tfrac{1}{16}&\tfrac{1}{4}&\tfrac{1}{4}&0\\ \tfrac{1}{16}&\tfrac{1}{4}&\tfrac{1}{4}&0\\ \tfrac{3}{8}&0&0&\tfrac{1}{2}\end{bmatrix}$.
This tensor does not correspond to the isotropic case. In isotropic linear viscoelasticity, the fourth-order viscosity tensor has the structure
@@ 5.1 Data generation
where $\alpha$ and $\theta$ are scalar constants. In such a case, certain components, such as $\mu_{1211}^{\rm iso}$, are identically zero. However, for the tensor defined above, we have $\mu_{1211}=\tfrac{1}{16}\neq 0$, thereby violating the isotropic form. This confirms that the chosen $\bm{\mu}$ represents an anisotropic medium. For simplicity, in this numerical section, we choose $\nabla p=0$ as the pressure gradient does not influence the formulation of our reconstruction method.
We employ a semi-implicit finite difference scheme to solve the forward problem numerically. The computational domain $\Omega=(-1,1)^{2}$ is discretized using a uniform $41\times 41$ Cartesian grid, corresponding to $\Delta x=\Delta y=0.05$. The temporal interval $[0,T]$ with $T=1$ is uniformly partitioned into 700 time steps $\{t^{1},t^{2},\dots,t^{700}\}$, with the step size $\Delta t\approx 0.00143$.
For the time discretization, we denote $\mathbf{u}(x,y,t^{n+1})$ by ${\bf u}^{n}$ and adopt a semi-implicit scheme in which the time derivative is approximated by $\frac{\mathbf{u}^{n+1}-\mathbf{u}^{n}}{\Delta t}.$ The diffusion term involving the fourth-order viscosity operator ${\rm div}(\bm{\mu}:\nabla\mathbf{u})$ is treated implicitly at time $t^{n+1}$, while the nonlinear convection term $(\mathbf{u}\cdot\nabla)\mathbf{u}$ is computed explicitly using the known velocity field $\mathbf{u}^{n}$ from the previous time step. This leads to a linear system for ${\bf u}^{n+1}$ at each time step, which is solved using a sparse direct solver. Homogeneous Dirichlet boundary conditions are imposed at each step. This implicit-explicit like approach improves stability while keeping the computational complexity manageable.
After obtaining the numerical solution $\mathbf{u}^{*}$ to the forward problem using the semi-implicit finite difference method described above, we extract the associated Neumann boundary data $\partial_{\nu}\mathbf{u}^{*}$ by computing the normal derivative of the velocity field along the boundary $\partial\Omega$. This exact boundary flux is then used to generate synthetic measurement data by introducing multiplicative noise. Specifically, we define the noisy observations $\mathbf{f}=\mathbf{f}^{*}(1+\delta\,\mathcal{R})$, where $\mathbf{f}^{*}=\partial_{\nu}\mathbf{u}^{*}$, $\delta=10\%$ is the noise level, and $\mathcal{R}$ is a uniformly distributed random field with values in [1,1][-1,1]. This formulation reflects realistic measurement errors and serves to validate the robustness of the proposed inversion algorithm under noisy boundary conditions.
@@ 5.2 Numerical examples
With the boundary data available, we proceed to solve the inverse problem. The key computational steps of our proposed method, based on the time dimension reduction framework, are outlined in Algorithm 1.
In Step 1 of Algorithm 1, we determine the set of parameters through a trial-and-error process, guided by the results of a reference test case. Specifically, we choose $N=30$ and $\epsilon=10^{-6}$. After fine-tuning these parameters to achieve optimal performance in the reference test, we use the same set of parameters for all subsequent tests to ensure consistency and comparability across different scenarios.
In Step 4, we compute ${\bf U}^{(k+1)}$ based on the current knowledge of ${\bf U}^{(k)}$. As stated earlier, ${\bf U}^{(k+1)}$ is the solution to the linear system in (4.1), which can be obtained through least-squares optimization. The minimization process, once discretized, can be performed using optimization packages available in Matlab. In our implementation, we use the lsqlin function, a built-in tool in Matlab, to solve the problem. For further details on the implementation of this Carleman-weighted least-squares optimization with lsqlin, we refer to [21, 26].
Other steps of Algorithm 1 can be implemented straightforwardly.
@@
In the first test, the true initial velocity field $\mathbf{U}_{0}^{\rm true}=\begin{bmatrix}u_{0}^{{\rm true},1}&u_{0}^{{\rm true},2}\end{bmatrix}^{\top}$ is constructed with high contrast between the regions of non-zero and background zero. Specifically, we set
$u_{0}^{{\rm true},1}(x,y)=\begin{cases}4,&\text{if }8(x-0.4)^{2}+y^{2}<0.64,\\ 0,&\text{otherwise},\end{cases}\quad u_{0}^{{\rm true},2}(x,y)=\begin{cases}4,&\text{if }x^{2}+8(y-0.3)^{2}<0.64,\\ 0,&\text{otherwise}.\end{cases}$
These profiles represent two elliptical regions: one vertically elongated in $u_{0}^{{\rm true},1}$ and the other horizontally elongated in $u_{0}^{{\rm true},2}$. This configuration allows us to test our method’s ability to recover both the magnitude and geometry of anisotropic features in the initial data, with high contrast between the active and inactive regions. The data is evolved forward under the anisotropic Navier–Stokes equations, and the resulting boundary observations are then used as input for the inverse reconstruction algorithm.
(a) $u_{0}^{{\rm true},1}$
(b) $u_{0}^{{\rm true},2}$
(c) $\ln\Big(\frac{\|{\bf U}^{(k)}-{\bf U}^{(k-1)}\|_{L^{\infty}}}{\|{\bf U}^{(k)}\|_{L^{\infty}}}\Big)$
(d) $u_{0}^{{\rm comp},1}$
(e) $u_{0}^{{\rm comp},2}$
Figure 1: Test 1: The true and computed solutions to the inverse initial data problem. The first row shows the true velocity fields and the logarithm of the error between consecutive iterations. The second row displays the computed velocity fields corresponding to the reconstruction.
The results presented in Figure 1 indicate a promising outcome for solving the inverse initial data problem, even in the presence of 10% noise in the data. The true velocity fields, shown in the first row, exhibit the expected localized features with distinct directional characteristics. The reconstructed velocity fields in the second row closely match the true data, demonstrating the effectiveness of the proposed method in accurately recovering the key structural elements of the original velocity field. From a quantitative perspective, the computed maximum value of $u_{0}^{{\rm comp},1}$ within the vertical ellipse is 4.1074 (with a relative error of 2.69%), and the computed maximum value of $u_{0}^{{\rm comp},2}$ within the horizontal ellipse is 4.0594 (with a relative error of 1.49%). Furthermore, the convergence plot (figure c) shows a consistent reduction in the error over successive iterations, emphasizing the method’s robustness. These results suggest that the method can provide reliable reconstructions, even with noisy observations, and holds potential for further refinement in more complex applications. Overall, the findings offer a positive outlook for the applicability of this approach in addressing real-world inverse problems effectively.
@@
For the second test case, the true initial velocity field consists of two components with localized features in the shape of diagonal ellipses. Specifically, the first component, $u_{0}^{\rm true,1}$, is defined as:
$u_{0}^{\rm true,1}(x,y)=\begin{cases}1,&\text{if }1.5(x+y)^{2}+y^{2}<0.7^{2},\\ 0,&\text{otherwise}.\end{cases}$
The second component, $u_{0}^{\rm true,2}$, takes the form:
$u_{0}^{\rm true,2}(x,y)=\begin{cases}1,&\text{if }x^{2}+1.5(x-y)^{2}<0.7^{2},\\ 0,&\text{otherwise}.\end{cases}$
@@
(a) $u_{0}^{{\rm true},1}$
(b) $u_{0}^{{\rm true},2}$
(c) $\ln\Big(\frac{\|{\bf U}^{(k)}-{\bf U}^{(k-1)}\|_{L^{\infty}}}{\|{\bf U}^{(k)}\|_{L^{\infty}}}\Big)$
(d) $u_{0}^{{\rm comp},1}$
(e) $u_{0}^{{\rm comp},2}$
Figure 2: Test 2: The true and computed solutions to the inverse initial data problem. The first row shows the true velocity fields and the logarithm of the error between consecutive iterations. The second row displays the computed velocity fields corresponding to the reconstruction.
The results displayed in Figure 2 demonstrate the effectiveness of the proposed method for solving the inverse initial data problem, even with 10% noise in the data. The true initial velocity fields, shown in the first row, exhibit diagonal elliptical structures with distinct orientations, accurately reflecting the expected features of the synthetic test case. In the second row, the computed velocity fields closely match the true data, highlighting the method’s ability to recover the key geometrical characteristics of the initial conditions despite the noisy observations. From a quantitative perspective, the computed maximum value of $u_{0}^{{\rm comp},1}$ within the ellipse is 0.8258 (with a relative error of 17.42%), and the computed maximum value of $u_{0}^{{\rm comp},2}$ within the ellipse is 0.8608 (with a relative error of 13.92%). These results show that the method is capable of capturing both the magnitude and shape of the velocity components, with errors remaining within an acceptable range. The logarithmic error plot in figure (c) illustrates the steady convergence of the method, providing further evidence of its robustness and accuracy in reconstructing the true data despite the presence of noise.
@@
In this test, the true initial velocity field is defined as follows. The first component, $u_{0}^{\rm true,1}(x,y)$, represents the characteristic function of an annular region. This is given by:
$u_{0}^{\rm true,1}(x,y)=\begin{cases}1,&\text{if }0.6^{2}<x^{2}+y^{2}<0.9^{2},\\ 0,&\text{otherwise}.\end{cases}$
The second component, $u_{0}^{\rm true,2}(x,y)$, is the characteristic function of a square with a central void, which is expressed as:
$u_{0}^{\rm true,2}(x,y)=\begin{cases}1,&\text{if }0.6<\max(|x|,|y|)<0.9,\\ 0,&\text{otherwise}.\end{cases}$
(a) $u_{0}^{{\rm true},1}$
(b) $u_{0}^{{\rm true},2}$
(c) $\ln\Big(\frac{\|{\bf U}^{(k)}-{\bf U}^{(k-1)}\|_{L^{\infty}}}{\|{\bf U}^{(k)}\|_{L^{\infty}}}\Big)$
(d) $u_{0}^{{\rm comp},1}$
(e) $u_{0}^{{\rm comp},2}$
Figure 3: Test 3: The true and computed solutions to the inverse initial data problem. The first row shows the true velocity fields and the logarithm of the error between consecutive iterations. The second row displays the computed velocity fields corresponding to the reconstruction.
The results displayed in Figure 3 highlight the effectiveness of our reconstruction method, even in the presence of 10% noise in the data. Despite the complexity of the true initial velocity field, which includes structures such as a ring-like formation and a square with a void, the reconstructed velocity fields (shown in the second row) closely match the true data (first row). This demonstrates the method’s ability to capture key structural features of the initial condition. The convergence plot in Figure (c) shows that the error between consecutive iterations consistently decreases, further emphasizing the robustness of the method. Quantitatively, the computed maximum value of $u_{0}^{{\rm comp},1}$ within the ring is 0.8756 (with a relative error of 12.44%), and the computed maximum value of $u_{0}^{{\rm comp},2}$ within the voided square is 0.9876 (with a relative error of 1.24%). These findings strongly suggest that the method holds significant promise for solving inverse initial data problems in real-world applications, even when the data is noisy.

### v4

@@ 5 Numerical study
In this section, we present numerical experiments to validate the proposed time-dimensional reduction method and to illustrate its performance in reconstructing the initial velocity field from noisy boundary observations.
5.1 Manufactured data generation
We present numerical simulations in the two-dimensional setting ($d=2$) for simplicity. In our experiments, the fourth-order viscosity tensor $\bm{\mu}$ is chosen to be spatially constant, and its flattened representation is given by
$\bm{\mu}_{\rm flat}=\begin{bmatrix}\mu_{1111}&\mu_{1112}&\mu_{1121}&\mu_{1122}\\ \mu_{1211}&\mu_{1212}&\mu_{1221}&\mu_{1222}\\ \mu_{2111}&\mu_{2112}&\mu_{2121}&\mu_{2122}\\ \mu_{2211}&\mu_{2212}&\mu_{2221}&\mu_{2222}\end{bmatrix}=\frac{1}{14}\begin{bmatrix}80&5&5&25\\ 5&30&5&0\\ 5&5&30&0\\ 25&0&0&40\end{bmatrix}$.
This tensor is anisotropic. Indeed, in the isotropic case, the fourth-order viscosity tensor must have the form
@@ 5.1 Manufactured data generation
where $\alpha$ and $\theta$ are scalar constants. In particular, for an isotropic tensor one has
$\mu_{1211}^{\rm iso}=0.$
However, for the tensor used here,
$\mu_{1211}=\frac{5}{14}\neq 0.$
Therefore, the above tensor does not satisfy the isotropic structure and hence represents an anisotropic medium.
To generate synthetic data, the density $\rho$, the pressure $p$ and the body force $\bf F$}, for the numerical experiments, we use a manufactured-solution approach. Let $\Omega=(-1,1)^{2}$ and $T=0.5$. Let ${\bf u}^{0}({\bf x})$ be the prescribed initial velocity field. We define the space-time velocity field by
${\bf u}({\bf x},t)=e^{-\lambda t}{\bf u}^{0}({\bf x})+(1-e^{-\lambda t}){\bf W}({\bf x}),\qquad{\bf x}=(x,y)\in\Omega,\ t\in(0,T),$ (5.1)
where $\lambda>0$ is fixed and
$\mathbf{W}(x,y)=\begin{pmatrix}(1-x^{2})(1-y^{2})\\[1.99997pt] (1-x^{2})(1-y^{2})\end{pmatrix}$. (5.2)
In our computational experiments, $\lambda=0.2$. Thus, ${\bf u}({\bf x},0)={\bf u}^{0}({\bf x})$, and the added profile $\bf W$} vanishes on $\partial\Omega$.
The density is initialized by a positive function $\rho_{0}({\bf x})=1$ and is then evolved numerically from the continuity equation
$\rho_{t}+{\rm div}(\rho{\bf u})=0\qquad\text{in }\Omega\times(0,T),$ (5.3)
using a finite-volume update with local Lax–Friedrichs numerical fluxes in both coordinate directions. In the experiments reported below, we set
$p({\bf x},t)=\rho({\bf x},t),$ (5.4)
which means $c$ in (1.1) is 1.
Once $\rho$ and $\bf u$} are available, the body force $\bf F$} is defined so that the momentum equation is satisfied numerically:
$\rho\bigl({\bf u}_{t}+({\bf u}\cdot\nabla){\bf u}\bigr)=-\nabla p+{\rm div}(\bm{\mu}:\nabla{\bf u})+\rho{\bf F}.$ (5.5)
More precisely, $\bf F$} is computed pointwise from
${\bf F}={\bf u}_{t}+({\bf u}\cdot\nabla){\bf u}+\frac{1}{\rho}\nabla p-\frac{1}{\rho}{\rm div}(\bm{\mu}:\nabla{\bf u}),$ (5.6)
where spatial derivatives are approximated by second-order finite differences, and time derivatives are approximated by central differences in the interior and one-sided differences at the endpoints. The anisotropic viscous term ${\rm div}(\bm{\mu}:\nabla{\bf u})$ is evaluated numerically using the prescribed fourth-order tensor $\bm{\mu}$. Therefore, the quadruple $(\rho,{\bf u},p,{\bf F})$ is consistent with the discrete forward model used to generate the boundary measurements.
After constructing the manufactured forward solution $(\rho,{\bf u},p,{\bf F})$ as described above, we extract the associated Neumann boundary data by computing the normal derivative $\partial_{\nu}{\bf u}$ along the boundary $\partial\Omega$. This exact boundary flux is then used to generate synthetic measurement data by introducing multiplicative noise. More precisely, we define the noisy observations by
$\mathbf{f}_{\rm noise}=\mathbf{f}(1+\delta\,\mathcal{R}),\qquad\mathbf{f}=\partial_{\nu}{\bf u},$
where $\delta=10\%$ is the noise level and $\mathcal{R}$ is a uniformly distributed random field taking values in [1,1][-1,1]. This construction is intended to mimic measurement errors and to assess the robustness of the proposed inversion algorithm under noisy boundary observations.
@@ 5.2 Numerical examples
With the boundary data available, we proceed to solve the inverse problem. The main computational steps of the proposed method, based on the time-dimensional reduction framework, are summarized in Algorithm 1.
In Step 1 of Algorithm 1, we select the truncation level $N$ and the regularization parameter $\epsilon$ through a trial-and-error calibration based on a reference test case. In the numerical experiments reported below, we choose
$N=20,\qquad\epsilon=10^{-8},$
and keep these values fixed across all test cases to ensure consistency and comparability across scenarios. We also use the damping parameter $\omega=0.5$ in the Picard iteration. The number of iteration $K$ in the loop Steps 47 of Algorithm 1 is 12.
In Step 5, the quasi-reversibility functional $J_{{\bf U}^{(k)}}^{\epsilon,N}$ is minimized to obtain an intermediate iterate $\widetilde{\bf U}^{(k+1)}$, which solves the linearized problem (4.1) in the least-squares sense. The new iterate ${\bf U}^{(k+1)}$ is then obtained through the damped Picard update introduced above. After spatial discretization, the minimization of $J_{{\bf U}^{(k)}}^{\epsilon,N}$ is carried out using standard optimization tools available in MATLAB. In our implementation, the spatial derivatives in (4.1) are discretized using finite differences on a uniform $41\times 41$ Cartesian grid, and the resulting least-squares problem is solved by the built-in MATLAB function lsqlin. For further details on the implementation of this least-squares optimization with lsqlin, we refer to [21, 26].
The remaining steps of Algorithm 1, including the reconstruction of the space-time approximation ${\bf u}^{\rm comp}$ and the recovery of the initial condition $\mathbf{U}_{0}^{\rm comp}$, are implemented directly from the definitions of the projected expansion.
@@
In the first test, the true initial velocity field ${\bf u}^{0}_{\rm true}=\begin{bmatrix}u^{0}_{{\rm true},1}&u^{0}_{{\rm true},2}\end{bmatrix}^{\top}$ is chosen so that its two components are supported in distinct elliptical regions. Specifically, we set
$u^{0}_{{\rm true},1}(x,y)=\begin{cases}1,&\text{if }8(x-0.4)^{2}+y^{2}<0.4^{2},\\ 0,&\text{otherwise},\end{cases}\qquad u^{0}_{{\rm true},2}(x,y)=\begin{cases}1,&\text{if }x^{2}+8(y-0.3)^{2}<0.4^{2},\\ 0,&\text{otherwise}.\end{cases}$
The function $u^{0}_{{\rm true},1}$ is supported in a vertically elongated ellipse centered near $(0.4,0)$, while $u^{0}_{{\rm true},2}$ is supported in a horizontally elongated ellipse centered near $(0,0.3)$. This choice allows us to examine the ability of the proposed method to recover both the location and the geometry of localized anisotropic structures in the initial data. The boundary observations are generated synthetically from the manufactured data $(\rho,{\bf u},p,{\bf F})$ and then used as input for the inverse reconstruction algorithm.
(a) True first velocity component $u^{0}_{\mathrm{true},1}$
(b) Reconstructed first velocity component $u^{0}_{\mathrm{comp},1}$
(c) Relative error of the first component $\frac{|u^{0}_{{\rm true},1}-u^{0}_{{\rm comp},1}|}{\|u^{0}_{{\rm true},1}\|_{L^{\infty}(\Omega)}}$
(d) True second velocity component $u^{0}_{\mathrm{true},2}$
(e) Reconstructed second velocity component $u^{0}_{\mathrm{comp},2}$
(f) Relative error of the second component $\frac{|u^{0}_{{\rm true},2}-u^{0}_{{\rm comp},1}|}{\|u^{0}_{{\rm true},2}\|_{L^{\infty}(\Omega)}}$
(g) True initial velocity field $\mathbf{u}^{0}_{\mathrm{true}}$
(h) Reconstructed initial velocity field $\mathbf{u}^{0}_{\mathrm{comp}}$
(i) Residual of (3.15) for ${\bf U}^{(k)}$ generated by Algorithm 1
Figure 1: Test 1: true and reconstructed initial velocity fields. Figures (a)–(b) compare the true and reconstructed first velocity component, while Figure (c) displays the corresponding relative error. Figures (d)–(e) compare the true and reconstructed second velocity component, while Figure (f) shows the corresponding relative error. Figures (g)–(h) present the true and reconstructed initial velocity fields, respectively. Figure (i) plots the residual of (3.15) for ${\bf U}^{(k)}$ generated by Algorithm 1, showing that ${\bf U}^{(k)}$ approximately solves the time-reduction model (3.15).
The results presented in Figure 1 demonstrate the strong performance of the proposed method for the inverse initial data problem, even in the presence of $10\%$ noise in the boundary measurements. Figures 1a1b and Figures 1d1e show that the reconstructed first and second velocity components recover not only the correct location and support of the true solution, but also its essential geometric structure. Although the reconstructions exhibit slight diffusion and mild background artifacts, the main features of both components are captured accurately. This is further confirmed by the relative error plots in Figures 1c and 1f, where the displayed maxima are approximately $0.035$ and $0.055$, respectively. In addition, the comparison between Figures 1g and 1h shows that the reconstructed initial velocity field preserves the dominant directional pattern of the true field remarkably well. Finally, Figure 1i reveals a clear and steady decay of the residual of (3.15) for ${\bf U}^{(k)}$, providing strong numerical evidence for the stability and effectiveness of the iterative reconstruction procedure. Altogether, these results highlight the robustness of the proposed approach and underscore its strong potential for solving noisy inverse problems in more realistic and challenging settings.
@@
For the second test case, the true initial velocity field comprises two components, each with localized features in the form of diagonal ellipses. Specifically, the first component, $u^{0}_{\rm true,1}$, is defined as:
$u^{0}_{\rm true,1}(x,y)=\begin{cases}1,&\text{if }1.5(x+y)^{2}+y^{2}<0.5^{2},\\ 0,&\text{otherwise}.\end{cases}$
The second component, $u^{0}_{\rm true,2}$, takes the form:
$u^{0}_{\rm true,2}(x,y)=\begin{cases}1,&\text{if }x^{2}+1.5(x-y)^{2}<0.5^{2},\\ 0,&\text{otherwise}.\end{cases}$
@@
(a) True first velocity component $u^{0}_{\mathrm{true},1}$
(b) Reconstructed first velocity component $u^{0}_{\mathrm{comp},1}$
(c) Relative error of the first component $\frac{|u^{0}_{{\rm true},1}-u^{0}_{{\rm comp},1}|}{\|u^{0}_{{\rm true},1}\|_{L^{\infty}(\Omega)}}$
(d) True second velocity component $u^{0}_{\mathrm{true},2}$
(e) Reconstructed second velocity component $u^{0}_{\mathrm{comp},2}$
(f) Relative error of the second component $\frac{|u^{0}_{{\rm true},2}-u^{0}_{{\rm comp},2}|}{\|u^{0}_{{\rm true},2}\|_{L^{\infty}(\Omega)}}$
(g) True initial velocity field $\mathbf{u}^{0}_{\mathrm{true}}$
(h) Reconstructed initial velocity field $\mathbf{u}^{0}_{\mathrm{comp}}$
(i) Residual of (3.15) for ${\bf U}^{(k)}$ generated by Algorithm 1
Figure 2: Test 2: true and reconstructed initial velocity fields. Figures (a)–(c) show the true first component, its reconstruction, and the corresponding relative error. Figures (d)–(f) show the true second component, its reconstruction, and the corresponding relative error. Figures (g)–(h) display the true and reconstructed initial velocity fields. Figure (i) shows the residual of (3.15) for ${\bf U}^{(k)}$ generated by Algorithm 1, which decreases steadily throughout the iteration.
The results shown in Figure 2 further demonstrate the effectiveness of the proposed method in reconstructing initial velocity fields with non-axis-aligned geometric structures. Figure 2a and Figure 2b, as well as Figure 2d and Figure 2e, indicate that both diagonal components are recovered with the correct orientation, location, and overall support. Although the reconstructed profiles are slightly smoothed and accompanied by mild background artifacts, the dominant diagonal features remain clearly visible. This is supported by the relative error distributions in Figure 2c and Figure 2f, whose displayed maxima are approximately $0.11$ and $0.07$, respectively. Moreover, the comparison between Figure 2g and Figure 2h shows that the reconstructed initial velocity field preserves the main directional structure of the true field. Finally, Figure 2i shows a steady decay of the residual of (3.15) for ${\bf U}^{(k)}$, which again provides numerical evidence for the stability and reliability of the damped Picard iteration. Overall, this test confirms that the proposed method is capable of accurately recovering localized anisotropic structures with different orientations, even under noisy boundary observations.
@@
In the third test, the true initial velocity field is chosen so that both components contain a multi-layer structure with an outer ring and an inner core of different magnitude. The first component, $u^{0}_{\rm true,1}(x,y)$, is defined by
$u^{0}_{\rm true,1}(x,y)=\begin{cases}1,&\text{if }0.6^{2}<x^{2}+y^{2}<0.7^{2},\\ 2,&\text{if }x^{2}+y^{2}<0.2^{2},\\ 0,&\text{otherwise},\end{cases}$
so that it consists of a circular outer annulus together with a central disk of larger positive magnitude. The second component, $u^{0}_{\rm true,2}(x,y)$, is defined by
$u^{0}_{\rm true,2}(x,y)=\begin{cases}1,&\text{if }0.6^{2}<\max(x^{2},y^{2})<0.7^{2},\\ -2,&\text{if }\max(x^{2},y^{2})<0.2^{2},\\ 0,&\text{otherwise},\end{cases}$
which represents a square outer ring together with a central square core of negative magnitude. This test is designed to assess whether the proposed method can recover composite structures with multiple active regions, sharp interfaces, and sign changes across different components of the initial velocity field.
(a) True first velocity component $u^{0}_{\mathrm{true},1}$
(b) Reconstructed first velocity component $u^{0}_{\mathrm{comp},1}$
(c) Relative error of the first component $\frac{|u^{0}_{{\rm true},1}-u^{0}_{{\rm comp},1}|}{\|u^{0}_{{\rm true},1}\|_{L^{\infty}(\Omega)}}$
(d) True second velocity component $u^{0}_{\mathrm{true},2}$
(e) Reconstructed second velocity component $u^{0}_{\mathrm{comp},2}$
(f) Relative error of the second component $\frac{|u^{0}_{{\rm true},2}-u^{0}_{{\rm comp},2}|}{\|u^{0}_{{\rm true},2}\|_{L^{\infty}(\Omega)}}$
(g) True initial velocity field $\mathbf{u}^{0}_{\mathrm{true}}$
(h) Reconstructed initial velocity field $\mathbf{u}^{0}_{\mathrm{comp}}$
(i) Residual of (3.15) for ${\bf U}^{(k)}$ generated by Algorithm 1
Figure 3: Test 3: true and reconstructed initial velocity fields. Figures (a)–(c) show the true first component, its reconstruction, and the corresponding relative error, respectively. Figures (d)–(f) show the true second component, its reconstruction, and the corresponding relative error, respectively. Figures (g)–(h) display the true and reconstructed initial velocity fields. Figure (i) shows the residual of (3.15) for ${\bf U}^{(k)}$ generated by Algorithm 1, which decreases steadily throughout the iteration.
The results shown in Figure 3 demonstrate that the proposed method remains effective for more intricate initial data containing multiple layers, disconnected active regions, and sign changes. Figure 3a and Figure 3b show that the first component is reconstructed with the correct radial organization: both the outer annular structure and the inner disk are clearly identified, with their locations and relative magnitudes reasonably preserved. Likewise, Figure 3d and Figure 3e indicate that the second component successfully recovers the square outer ring together with the negative inner core, which confirms that the method is capable of distinguishing structures of different geometry and opposite sign within the same test. Although the reconstructed profiles exhibit noticeable smoothing and mild background artifacts, the essential topology and contrast pattern of the true solution are retained. This is further supported by the relative error distributions in Figure 3c and Figure 3f, where the displayed maxima are approximately $0.13$ and $0.14$, respectively. In addition, the comparison between Figure 3g and Figure 3h shows that the reconstructed initial velocity field captures the main directional organization of the true field. Finally, Figure 3i shows a steady decay of the residual of (3.15) for ${\bf U}^{(k)}$, providing further numerical evidence for the stability and robustness of the damped Picard iteration. Overall, this example illustrates that the proposed method can successfully recover composite structures with multiple scales, sharp interfaces, and sign-changing behavior from noisy boundary observations.